OpenStudy (kirbykirby):

An iid sample \(X_1,\dots\,X_n\) have pdf \(f(x;\theta)=\theta x^{\theta-1},\,\,\,\, 00\). The MLE is \[\hat{\theta}=\frac{-n}{\sum_{i=1}^{n}\log(X_i)}\]. I need to find the expectation of \(\hat{\theta}\).

OpenStudy (kirbykirby):

It suggests showing \(-\log X_i\)~EXP\((1,\theta)\), then \(-\sum_{i=1}^n \log X_i\)~GAM\((n, 1/\theta)\), which I did and is ok. The next step though says \[E\left[\left(-\sum_{i=1}^n \log X_i\right)^{-1}\right]=\left(\frac{1}{\theta}\right)^{-1}\frac{\Gamma(n-1)}{\Gamma(n)}=\frac{\theta}{n-1}\] I understand that the expected value for GAM\((n, 1/\theta)\) is \(n/\theta\), so I'm not sure how they got the line above??

OpenStudy (kirbykirby):

Do we use another 1-1 transformation using Y = 1/X, and then calculate the expectation using the definition, or is there a faster way to do this?